• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 12:53:22 on October 27,1998

    Entry number 3839

    keyword=Hall A harp Scan

    We made a harp scan of the last harp #6 close to BPM1H03B ala Pascal's instructions. The data is recorded in file: oct2798-6-1. The beam current was 10 microA cw.


    FIGURE 1